# Extension principles for the Einstein Yang–Mills system

### v1

@@ Appendix A Local existence theorem
[…]
@@
[…]
Convergence of $\{\hat{w}_{k}\}$ uniformly in the energy norm. By (A.13) and the fact $\varepsilon$ is small, the estimate (A) tells that
@@
[…]
Moreover, by virtue of the fact $\frac{|\hat{w}_{k+1}-\hat{w}_{k}|^{2}}{r_{k-1}^{2}}\lesssim\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{1}(C_{u})}$, the previous $L^{\infty}$ estimates and (A.13), it follows that
@@
$\frac{m_{k+1}}{r_{k}}-\frac{m_{k}}{r_{k-1}}={}$ $\frac{m_{k+1}-m_{k}}{r_{k}}+\frac{m_{k}}{r_{k}}\left(1-\frac{r_{k}}{r_{k-1}}\right),$ (A.19)
and using (A.3), we obtain
@@
Then it follows from the inductive assumptions (A.13) and (A) that
$|\mu_{k+1}-\mu_{k}|\leq\varepsilon CM2^{-k+1}.$ (A.20)
@@
Furthermore, in view of the formulae (3.1.6) and (3.1.6), it follows from (A) and (A.20) that (A.14)–(A.15) hold with $k$ replaced by $k+1$.
Convergence of $\Omega^{2}_{k}h_{k}$, $\underline{h}_{k}$, $h_{k}$, $m_{k}$ (uniformly in $C^{0}$ norm). Similarly, using the equations (A.5) and (A.6), we prove that
@@
which together with (A.20) implies $|h_{k+1}^{-1}-h^{-1}_{k}|\leq\varepsilon M2^{-k}$. The inductive proof is thus complete.
Finally, substituting all the previous results into the equation of mass (A.3), we obtain
@@
Along with the formula of $\mu$, $m$ and $r$, the relation $h_{k}=\frac{\mu_{k}-1}{\underline{h}_{k}}$ gives rise to an identity
@@
we conclude $(r,\,m,\,w)$ satisfies the SSEYM equations. However, since $\partial_{u}\partial_{r}w$ is not necessarily bounded on the axis, the Yang–Mills equation (and also the whole SSEYM system) holds in the sense of distribution.
Continuous dependence on data. Suppose $\hat{w}_{n}$ and $\hat{w}_{m}$ are solutions with data $\hat{w}_{0n}$, $\hat{w}_{0m}$ respectively. The following equation holds
@@
Appendix B An extension principle away from the axis
The extension principle away from the axis had been confirmed for systems, such as Einstein-Vlasov and Einstein charged scalar field in spherical symmetry [14, 15]. A prior to an extension theorem, a local existence theorem away from the axis is needed and it was addressed in [14] as well. The proof in [14] was based on the fixed point method and the double null gauge. This scheme indeed works well when the Einstein equations are coupled with other matter fields. In this section, we follow the approaches in [14, 15] to develop the local theory associated to the SSEYM system.
@@ Appendix B An extension principle away from the axis
$r\partial_{u}\partial_{\underline{u}}r={}$ $-\Omega^{2}-\partial_{u}r\partial_{\underline{u}}r+\Omega^{2}r^{-2}Q^{2},$ (B.1a)
$r^{2}\partial_{u}\partial_{\underline{u}}\ln\Omega={}$ $\Omega^{2}+\partial_{u}r\partial_{\underline{u}}r-2\Omega^{2}r^{-2}Q^{2},$ (B.1b)
$\partial_{u}\left(\Omega^{-2}\partial_{u}r\right)={}$ $-2\Omega^{-2}r^{-2}\left(\partial_{u}w\right)^{2},$ (B.1c)
$\partial_{\underline{u}}\left(\Omega^{-2}\partial_{\underline{u}}r\right)={}$ $-2\Omega^{-2}r^{-2}\left(\partial_{\underline{u}}w\right)^{2},$ (B.1d)
$\partial_{u}\partial_{\underline{u}}w+{}$ $\Omega^{2}r^{-2}wQ=0,$ (B.1e)
where $Q=w^{2}-1$. In view of the structures of the system (B.1a)-(B.1e), a local existence theorem away from the axis follows in a standard way [14].
Theorem B.1.
Let $k\geq 0$ and $\Sigma:=[0,d]\times\{0\}\cup\{0\}\times[0,d]$. On $\Sigma$, let $r,\Omega$ be positive functions such that $r\in C^{k+2}(\Sigma)$ and $\Omega\in C^{k+1}(\Sigma)$, and let $w\in C^{k+1}(\Sigma)$. Suppose (B.1c)–(B.1d) hold initially on $[0,d]\times\{0\}$ and $\{0\}\times[0,d]$, respectively. Define
@@
Then there exists a $\delta>0$ depending only on $N$ such that there are functions $(r,\Omega,w)\in C^{k+2}\times C^{k+1}\times C^{k+1}$ satisfying equations (B.1a)–(B.1e) in $[0,\delta^{\ast}]\times[0,\delta^{\ast}]$, where $\delta^{\ast}=\min\{d,\,\delta\}$, with the above prescribed data on $\Sigma$.
With respect to the above local theorem, we can prove a generalized extension principle away from the axis (i.e. “strongly tame” in the language of [15]) following [15].
Theorem B.2.
@@
The proof relies on the dominant energy condition and the special structure of Einstein Yang–Mills system. This structure manifests itself in that to bound all the variables, it suffices to bound
@@
which in the SSEYM case takes the form of
@@
Nevertheless, this quantity can be bounded straightforwardly through integrating (B.1a), please refer to [15] for more details.
Theorem B.2 shows that the “first singularity” in the maximal future development of spherically symmetric asymptotically flat data for the SSEYM system must emanate from the axis.
Appendix C Density theorem
As explained in Remark 1.6 (or Remark 1.2), regarding to the energy spaces in the extension theorems, we consider their equivalence, the weighted Sobolev space $H^{\prime s}_{w}$ with finite $H^{\prime s}_{w}$ norm [^4] If we take $r_{\ast}=0$ and $s=2$, the $H^{\prime s}_{w}$ space reduces to the $H^{1}$ space (1.5). , where
$\|\phi\|_{H^{\prime s}_{w}}:=\sum_{k\leq s}\int_{\{r_{\ast}<r\}}\xi^{-3-\gamma}r^{-1-b}\left((\xi r\partial_{r})^{k}\phi\right)^{2}\mathrm{d}r,\quad s\in\mathbb{Z}_{\geq 0}.$
We denote
$\rho:=\xi^{-1}r^{-1},\quad\text{that is}\,\,\rho=(r-r_{\ast})^{-1}.$
If $r_{\ast}=0$, $\xi=1$, and
$\xi^{-3-\gamma}r^{-1-b}=\rho^{1+b};$
If $r_{\ast}>0$, there is
$\xi^{-3-\gamma}r^{-1-b}\sim\rho^{3+\gamma}+\rho^{1+b}.$
Thus the $H^{\prime s}_{w}$ space is equivalent to the Sobolev space $H^{s}_{w}$ which has finite $H^{s}_{w}$ norm with
$\|\phi\|^{2}_{H^{s}_{w}}:=\sum_{k\leq s}\int_{\{r_{\ast}<r\}}f(\rho)\left((\rho^{-1}\partial_{r})^{k}\phi\right)^{2}\mathrm{d}r.$
Here $f(\rho)=(\rho^{3+\gamma}+\rho^{1+b})$ if $r_{\ast}>0$, and $f(\rho)=\rho^{1+b}$ if $r_{\ast}=0$. One can refer to [22, Section 3.2.4.] for more details of the $H^{s}_{w}$ space. In the remainder of this section, we aim to establish the density theorem in the weighted Sobolev space $H^{s}_{w}$.
Theorem C.1.
Smooth functions with compact support in $(r_{\ast},+\infty)$ is dense in $H^{s}_{w}\left((r_{\ast},+\infty)\right)$, $s\in\mathbb{Z}_{\geq 0}$.
Remark C.2.
If we take $r_{\ast}=0$, Theorem C.1 amounts to stating the density theorem in the $H^{1}$ space (1.5).
Proof.
The domain $(r_{\ast},+\infty)$ has an open covering
$(r_{\ast},+\infty)=\cup_{i\geq 0}\mathcal{D}_{1i},$
where we define
$\mathcal{D}_{1i}:=\{x\in(r_{\ast},+\infty)|\quad 2^{i}<\rho(x)<2^{i+3}\},\quad i\geq 1,$
and
$\mathcal{D}_{10}:=\{x\in(r_{\ast},+\infty)|\quad\rho(x)<2^{3}\}.$
On the other hand, letting
$\mathcal{D}_{2i}:=\{x\in(r_{\ast},+\infty)|\quad 2^{-i+1}<\rho(x)<2^{-i+1+3}\},\quad i\geq 1,$
we have
$\mathcal{D}_{10}\subset\cup_{i=1}^{+\infty}\mathcal{D}_{2i}.$
Therefore, $(r_{\ast},+\infty)$ is also covered by
$(r_{\ast},+\infty)=\cup_{i\geq 1}\left(\mathcal{D}_{1i}\cup\mathcal{D}_{2i}\right).$
Suppose the open covering $\cup_{i\geq 1}\left(\mathcal{D}_{1i}\cup\mathcal{D}_{2i}\right)$ which is locally finite, is associated to the partition of unity $\{\psi_{1i},\,\psi_{2i}\}_{i\geq 1}$. Namely,
$0\leq\psi_{ki}\leq 1,\,\psi_{ki}\in C_{0}^{\infty}(\mathcal{D}_{ki}),\,\,k=1,2,$
$\sum_{i\geq 1}\psi_{1i}+\psi_{2i}=1,$
and for all $m\in\mathbb{Z}_{>0}$,
[…]
The existence of such partition of unity is proved as follows. Note that, $\mathcal{D}_{1i}$ and $\mathcal{D}_{2i}$ can be alternatively written as
$\mathcal{D}_{1i}:=\{r\in(r_{\ast},+\infty)|\quad 2^{-i-3}<r-r_{\ast}<2^{-i}\},$
and
$\mathcal{D}_{2i}:=\{r\in(r_{\ast},+\infty)|\quad 2^{i-1-3}<r-r_{\ast}<2^{i-1}\},$
for $i\geq 1$. Take a smooth function $\phi$ such that
$\text{supp}\,\phi\subset\left(2^{-3},\,1\right).$
Then we define
$\phi_{1i}:=\phi\left(2^{i}(r-r_{\ast})\right),\quad\phi_{2i}:=\phi\left(2^{-i+1}(r-r_{\ast})\right)$
and let
$\psi_{1i}=\frac{\phi_{1i}}{\sum_{i\geq 1}\phi_{1i}+\phi_{2i}},\quad\psi_{2i}=\frac{\phi_{2i}}{\sum_{i\geq 1}\phi_{1i}+\phi_{2i}}.$
Now $\{\psi_{1i},\,\psi_{2i}\}_{i\geq 1}$ is the required partition of unity.
We can approximate $f\in H^{s}_{w}$ by the sequence $\sum_{j=1}^{K}\left(\psi_{1j}+\psi_{2j}\right)f$, $K\geq 1$, so that
$\|f-\sum_{j=1}^{K}\left(\psi_{1j}+\psi_{2j}\right)f\|_{H^{s}_{w}}\rightarrow 0,\quad\text{if}\,\,K\rightarrow+\infty.$
Furthermore, each of $\psi_{1j}f,\,\psi_{2j}f$, $1\leq i\leq K$, which belongs to the standard Sobolev space $H^{s}(\mathcal{D}_{1j})$, $H^{s}(\mathcal{D}_{2j})$ respectively can be approximated by $C_{0}^{\infty}$ functions. We thus conclude $C_{0}^{\infty}\left((r_{\ast},\,+\infty)\right)$ is dense in $H^{s}_{w}\left((r_{\ast},\,+\infty)\right)$. ∎

### v2

@@ 4. Local existence theorem
$:={}$ $\int_{C_{u}}\left(|\partial_{r_{k-1}}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)|^{2}+\frac{|\hat{w}_{k+1}-\hat{w}_{k}|^{2}}{r_{k-1}^{4}}\right)\Omega_{k-1}^{2}h_{k-1}\mathrm{d}\underline{u},$
@@
$:={}$ $\int_{C_{u}}|r_{k-1}\partial^{2}_{r_{k-1}}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)|^{2}\Omega_{k-1}^{2}h_{k-1}\mathrm{d}\underline{u}+\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{1}(C_{u})}.$
Convergence of $\{\hat{w}_{k}\}$ uniformly in the energy norm. By (4.13) and the fact $\varepsilon$ is small, the estimate (4) tells that
@@
$\leq{}$ $\varepsilon CM^{2}2^{-2k+2}.$ (4.17)
Moreover, by virtue of the fact $\frac{|\hat{w}_{k+1}-\hat{w}_{k}|^{2}}{r_{k-1}^{2}}\lesssim\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{1}(C_{u})}$, the previous $L^{\infty}$ estimates and (4.13), it follows that
@@
$\frac{m_{k+1}}{r_{k}}-\frac{m_{k}}{r_{k-1}}={}$ $\frac{m_{k+1}-m_{k}}{r_{k}}+\frac{m_{k}}{r_{k}}\left(1-\frac{r_{k}}{r_{k-1}}\right),$ (4.19)
and using (4.3), we obtain
@@
Then it follows from the inductive assumptions (4.13) and (4) that
$|\mu_{k+1}-\mu_{k}|\leq\varepsilon CM2^{-k+1}.$ (4.20)
@@
Furthermore, in view of the formulae (3.7) and (3.7), it follows from (4) and (4.20) that (4.14)–(4.15) hold with $k$ replaced by $k+1$.
Convergence of $\Omega^{2}_{k}h_{k}$, $\underline{h}_{k}$, $h_{k}$, $m_{k}$ (uniformly in $C^{0}$ norm). Similarly, using the equations (4.5) and (4.6), we prove that
@@
which together with (4.20) implies $|h_{k+1}^{-1}-h^{-1}_{k}|\leq\varepsilon M2^{-k}$. The inductive proof is thus complete.
Finally, substituting all the previous results into the equation of mass (4.3), we obtain
@@
Besides, the identity $h_{k}=\frac{\mu_{k}-1}{\underline{h}_{k}}$ gives rise to
@@
we conclude that (r,m,w)(r,\,m,\,w) satisfies the SSEYM equations. However, since $\partial_{u}\partial_{r}w$ is not necessarily bounded on the axis, the Yang–Mills equation, and indeed the entire SSEYM system, holds in the sense of distribution.
Continuous dependence on the data. Suppose $\hat{w}_{n}$ and $\hat{w}_{m}$ are solutions with data $\hat{w}_{0n}$, $\hat{w}_{0m}$ respectively. The following equation holds
@@
Appendix A An extension principle away from the axis
The extension principle away from the axis had been established for several systems, including the spherically symmetric Einstein-Vlasov and Einstein charged scalar field models [14, 15]. As a prerequisite for such an extension theorem, a local existence theorem away from the axis is needed, which was also provided in [14] as well. The proof in [14] was based on the fixed point method and the double null gauge. This approach remains valid for the Einstein equations coupled with various matter fields. In this section, we adapt the methods of [14, 15] to develop a local theory to the SSEYM system away from the axis.
@@ Appendix A An extension principle away from the axis
$r\partial_{u}\partial_{\underline{u}}r={}$ $-\Omega^{2}-\partial_{u}r\partial_{\underline{u}}r+\Omega^{2}r^{-2}Q^{2},$ (A.1a)
$r^{2}\partial_{u}\partial_{\underline{u}}\ln\Omega={}$ $\Omega^{2}+\partial_{u}r\partial_{\underline{u}}r-2\Omega^{2}r^{-2}Q^{2},$ (A.1b)
$\partial_{u}\left(\Omega^{-2}\partial_{u}r\right)={}$ $-2\Omega^{-2}r^{-2}\left(\partial_{u}w\right)^{2},$ (A.1c)
$\partial_{\underline{u}}\left(\Omega^{-2}\partial_{\underline{u}}r\right)={}$ $-2\Omega^{-2}r^{-2}\left(\partial_{\underline{u}}w\right)^{2},$ (A.1d)
$\partial_{u}\partial_{\underline{u}}w+{}$ $\Omega^{2}r^{-2}wQ=0,$ (A.1e)
where $Q=w^{2}-1$. In view of the structures of the system (A.1a)-(A.1e), a local existence theorem away from the axis follows in a standard way [14].
Theorem A.1.
Let $k\geq 0$ and $\Sigma:=[0,d]\times\{0\}\cup\{0\}\times[0,d]$. On $\Sigma$, let $r,\Omega$ be positive functions such that $r\in C^{k+2}(\Sigma)$ and $\Omega\in C^{k+1}(\Sigma)$, and let $w\in C^{k+1}(\Sigma)$. Suppose (A.1c)–(A.1d) hold initially on $0,d]\times\{0\}$ and $\{0\}\times[0,d]$, respectively. Define
@@
Then there exists a $\delta>0$ depending only on NN such that there are functions $(r,\Omega,w)\in C^{k+2}\times C^{k+1}\times C^{k+1}$ satisfying equations (A.1a)–(A.1e) in $0,\delta^{\ast}]\times[0,\delta^{\ast}]$, where $\delta^{\ast}=\min\{d,\,\delta\}$, with the above prescribed data on $\Sigma$.
Building upon the local existence theorem established above, we can prove a generalized extension principle away from the axis (referred to as “strongly tame” in the terminology of [15]) by adapting the methods of [15].
Theorem A.2.
@@
The proof relies on the dominant energy condition and the special structure of the Einstein Yang–Mills system. This structure is reflected in the fact that bounding all variables reduces to estimating the integrals
@@
which in the SSEYM case takes the form
@@
This quantity can be bounded straightforwardly through integrating (A.1a); for further details, we refer to [15].
Theorem A.2 shows that the “first singularity” in the maximal future development of spherically symmetric asymptotically flat data for the SSEYM system must emanate from the axis.